# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ 3.2 Estimation of the Optimal Solution
We give a method for estimating the optimal solution ${\bm{x}}^{\ast}$ at each $t\geq 1$. Recall that, in this paper, we cannot observe the value of $F({\bm{x}})$, which is the optimization target, but only observe the value of $f({\bm{x}},{\bm{w}})$ with noise. As a result, we cannot estimate ${\bm{x}}^{\ast}$ directly from the observed value. Therefore, based on the estimate of $F({\bm{x}})$ calculated from the posterior mean of $f$, we define the estimated solution of ${\bm{x}}^{\ast}$, $\hat{\bm{x}}_{t}$, as follows:
$\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}})).$ (3.3)
Table 3: The values of ${\rm lcb}_{t}({\bm{x}})$ and ${\rm ucb}_{t}({\bm{x}})$ for commonly used robustness measures (Table 3 in Inatsu et al., 2024a )
@@
$\mathcal{M}(\cdot)$: Monotonic Lipschitz continuous map with a Lipschitz constant $K$
@@
The expectation and probability are taken with respect to the distribution of ${\bm{w}}$, and $\alpha$-value-at-risk is the same meaning as $\alpha$-quantile.
@@ 3.3 Acquisition Function
We give acquisition functions to determine the next point $({\bm{x}}_{t},{\bm{w}}_{t})$ to be evaluated. In this study, the estimated solution $\hat{\bm{x}}_{t}$ and the next point ${\bm{x}}_{t}$ to be evaluated are not necessarily the same. In previous studies (Kirschner et al.,, 2020; Inatsu et al., 2024a, ) on BO for robustness measures under the presence of environmental variables, ${\bm{x}}_{t}$ is determined based on the upper bound of the credible interval of a robustness measure, and ${\bm{w}}_{t}$ is determined based on the maximization of the posterior variance of $f({\bm{x}}_{t},{\bm{w}})$. We modify the method for selecting ${\bm{x}}_{t}$ and partially adopt the above approach. Since ${\bm{w}}_{t}$ is selected based on the posterior variance of $f$, this has the advantage that there are no hyper-parameters in acquisition functions, whereas the credible interval of $F({\bm{x}})$ depends on the hyper-parameter $\beta_{t}$. Therefore, we avoid explicitly determining $\beta_{t}$ and adopt to use $\beta_{t}$ as a realization from a probability distribution. For example, Takeno et al., (2023) proposed IRGP-UCB in the usual BO framework, where $\beta_{t}$ of GP-UCB is a random sample from a two-parameter exponential distribution, and Inatsu et al., 2024b proposed randomized straddle in the level set estimation framework, where the parameter $\beta_{t}$ of the straddle acquisition function (Bryan et al.,, 2005) is a random sample from a chi-square distribution. These methods not only do not require explicit specification of hyper-parameters, but also provide tighter theoretical bounds than ordinary GP-UCB and straddle. However, the target function of the problem setting above is $f$, and the theoretical guarantee of their algorithms strongly depends on $f$ following GPs. In contrast, $F({\bm{x}})$, which is dealt with in this paper, does not generally follow GPs, even if $f$ follows a GP. Therefore, to derive the theoretically valid method, we modify the method of selecting ${\bm{x}}_{t}$ based on the upper bound of the credible interval of $F({\bm{x}})$. Before explaining this, we first extract an equivalent representation of the standard GP-UCB and focus only on essential properties of GP-UCB. The following lemma holds for GP-UCB in single-objective optimization without input uncertainty.
@@
For a black-box function $f({\bm{x}})$ following GP, let $\mu_{t-1}({\bm{x}})$ be a posterior mean, $\sigma^{2}_{t-1}({\bm{x}})$ be a posterior variance, and $\beta_{t}\geq 0$ be a parameter. Define ${\rm u}_{t-1}({\bm{x}})=\mu_{t-1}({\bm{x}})+\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}})$, ${\rm l}_{t-1}({\bm{x}})=\mu_{t-1}({\bm{x}})-\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}})$ and
${\bm{x}}^{({\rm u})}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}{\rm u}_{t-1}({\bm{x}}),$
$\tilde{\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm u}_{t-1}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm l}_{t-1}({\bm{x}}))_{+},$
$\check{\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\mu_{t-1}({\bm{x}}),$
${\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\{\tilde{\bm{x}}_{t},\check{\bm{x}}_{t}\}}({\rm u}_{t-1}({\bm{x}})-{\rm l}_{t-1}({\bm{x}})),$
where $(a)_{+}$ is $a$ if $a>0$, and otherwise 0. Then, the equality ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}({\bm{x}}_{t})$ holds.
The proof is given in Appendix C. By using this lemma, we select ${\bm{x}}_{t}$ as follows.
@@
Let $\xi_{1},\ldots,\xi_{t}$ be random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Suppose that ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ are respectively lower and upper bounds satisfying (3.2), and $\hat{\bm{x}}_{t}$ is given by (3.3). Then, ${\bm{x}}_{t}$ is selected as follows:
@@
where $\tilde{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm ucb}_{t}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t}({\bm{x}}))_{+}$.
Finally, based on ${\bm{x}}_{t}$ obtained by (3.4), we select ${\bm{w}}_{t}$ as follows.
@@
The next environmental variable to be evaluated, ${\bm{w}}_{t}$, is selected as follows:
@@ 4 Theoretical Analysis
In this section, we provide theoretical guarantees for expected regrets in the proposed algorithm. Details of the proof are given in Appendix C. First, in order to evaluate the goodness of the estimated solution, we define the regret $r_{t}$ and cumulative regret $R_{t}$ as follows:
@@
In addition, to derive the theoretical guarantee in the proposed method, we introduce the maximum information gain $\gamma_{t}$. The maximum information gain is a measure often used in theoretical analysis of the GP-based BO and level set estimation (Srinivas et al.,, 2010; Bogunovic et al.,, 2016; Gotovos et al.,, 2013) and can be expressed as follows:
$\gamma_{t}=\frac{1}{2}\sup_{\{(\tilde{\bm{x}}_{1},\tilde{\bm{w}}_{1}),\ldots,(\tilde{\bm{x}}_{t},\tilde{\bm{w}}_{t})\}\subset\mathcal{X}\times\Omega}\log{\rm det}({\bm{I}}_{t}+\sigma^{-2}_{{\rm noise}}\tilde{\bm{K}}_{t}),$
where $\tilde{\bm{K}}_{t}$ is the $t\times t$ whose $(i,j)$-th element is $k((\tilde{\bm{x}}_{i},\tilde{\bm{w}}_{i}),(\tilde{\bm{x}}_{j},\tilde{\bm{w}}_{j}))$. For example, the order of $\gamma_{t}$ for commonly used kernel functions such as linear, Gaussian and Matérn kernels is sublinear under mild conditions (see, e.g., Theorem 5 in Srinivas et al., (2010)). Let $h(a):[0,\infty)\to[0,\infty)$ be a non-decreasing and concave function with $h(0)=0$, and $\mathcal{H}$ be a set of all $h(a)$. Then, we define the set of functions $q(a):[0,\infty)\to[0,\infty)$, $\mathcal{Q}$, as follows:
@@
For $\mathcal{Q}$, ${\rm ucb}_{t-1}({\bm{x}}_{t})$, ${\rm lcb}_{t-1}({\bm{x}}_{t})$ and $2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})$, we introduce the following assumption.
@@
There exists a function $q(x)\in\mathcal{Q}$ such that for any $t\geq 1,{\bm{x}}_{t},\beta_{t}$ and $\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})$, the following inequality holds:
${\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})\leq q(2\beta^{1/2}_{t}\max_{{\bm{w}}\in\Omega}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})).$
From Inatsu et al., 2024a , $q(x)$ for commonly used robustness measures including the expectation measure satisfies Assumption 4.1. We give representative robustness measures and their $q(x)$ in Table 4.333Inatsu et al., 2024a derives $q(a)$ for the standard deviation and variance, but uses a constant related to the norm on the reproducing kernel Hilbert space, which is different from the settings in this paper and is not included. Also, the probabilistic threshold is not included because $q(a)$ is not shown. Regarding the distributionally robust, monotonic Lipschitz map and weighted sum, $q(a)\in\mathcal{Q}$ is a non-decreasing function with $q(0)=0$, and for any $\alpha_{1},\alpha_{2}\geq 0$ and $q_{1}(a),q_{2}(a)\in\mathcal{Q}$, it is obtained from $\alpha_{1}q_{1}(a),\alpha_{1}q_{1}(a)+\alpha_{2}q_{2}(a)\in\mathcal{Q}$.
Table 4: Specific forms of $q(a)$ for commonly used robustness measures (modified version of Table 4 in Inatsu et al., 2024a )
@@
$q(a;F)$: a function $q(a)$ for $F({\bm{x}})$ satisfying Assumption 4.1
$\mathcal{M}(\cdot)$: Monotonic Lipschitz continuous map with a Lipschitz constant $K$
$q^{(\cdot)}(a)$: a function $q(a)$ for $F^{(\cdot)}({\bm{x}})$ satisfying Assumption 4.1
@@
The expectation and probability are taken with respect to the distribution of ${\bm{w}}$,
and $\alpha$-value-at-risk is the same meaning as $\alpha$-quantile
@@
Assume that (3.2), Regularity assumption and Assumption 4.1 hold. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Let $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be a function satisfying Assumption 4.1. Then, the following holds if Algorithm 1 is performed:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij}}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
If $\gamma_{t}$ is a sublinear function, the following equality holds:
@@
Then, if $h_{i}(a)$ is continuous at 0, from $h_{i}(0)=0$ and Theorem 4.1, $\mathbb{E}[R_{t}]/t$ satisfies
@@
As a special case, from Table 4, if the target robustness measure is the expectation measure, $q(a)=a$, that is, $n=\zeta_{i}=s_{i}=\lambda_{ij}=\nu_{ij}=1$ and $h_{1}(a)=a$. In this case, $\mathbb{E}[R_{t}]$ satisfies
@@
For six commonly used robustness measures including the expectation measure, the following corollary holds.
@@
Under the assumption in Theorem 4.1, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following inequality holds:
@@
where $C_{0}=2(2\log(|\mathcal{X}\times\Omega|)+2)/\log(1+\sigma^{-2}_{{\rm noise}})$, and $C$ is 8 for the mean absolute deviation measure, and 4 for other measures.
While Theorem 4.1 and (4.2) provide a guarantee on the expected value of the cumulative regret, they do not directly provide a guarantee on the expected value of the regret. Specifically, they do not answer the question of which estimated solution $\hat{\bm{x}}_{i}$, $1\leq i\leq t$ has the smallest $\mathbb{E}[r_{i}]$. Therefore, we define the index $\hat{t}$ of the optimal estimated solution up to time $t$ as follows:
$\hat{t}=\operatornamewithlimits{argmin}_{1\leq i\leq t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})].$ (4.3)
Here, $\mathbb{E}_{t-1}[\cdot]$ is the conditional expectation given $D_{t-1}$, where
@@
Under the assumption in Theorem 4.1, the following inequality holds:
@@
where $\hat{t}$ is given by (4.3), and $h_{i}(\cdot)$ and each coefficient are given in Theorem 4.1. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
@@
From Theorem 4.2, if $\gamma_{t}$ is a sublinear function and $h_{i}(\cdot)$ is continuous at 0, since $h_{i}(0)=0$ and (4.1), $\mathbb{E}[r_{\hat{t}}]$ satisfies
@@
Therefore, the expected regret at index $\hat{t}$ converges to 0. However, to obtain $\hat{t}$, we need to solve (4.3), which cannot be solved analytically in general. Nevertheless, $\hat{t}$ can be written as
@@
Furthermore, the posterior distribution of $f$ given $D_{t-1}$ follows GPs. By using this, we can generate sample paths $\hat{f}_{1},\ldots,\hat{f}_{M}$ of $f$ from the posterior distribution, calculate the value of $F({\bm{x}})$ for each $\hat{f}_{j}$, and estimate $\mathbb{E}_{t-1}[F({\bm{x}})]$ from these $M$ values. Therefore, we can estimate $\hat{t}$ from the sample paths. On the other hand, for the expectation measure, the following theorem shows that we can use $t$ as $\hat{t}$.
@@
Under the assumption in Theorem 4.1, the following holds for the expectation measure:

### v2

@@ 3.2 Estimation of Optimal Solution
We present a method for estimating the optimal solution ${\bm{x}}^{\ast}$ at each iteration $t\geq 1$. Recall that in this setting, the value of the objective function $F({\bm{x}})$ is unobservable; we can only access noisy observations of $f({\bm{x}},{\bm{w}})$. As a result, directly estimating ${\bm{x}}^{\ast}$ from the observed data is not possible. To address this, we define the estimated solution $\hat{\bm{x}}_{t}$ based on the estimate of $F({\bm{x}})$ calculated from the posterior mean of $f$ as follows:
$\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho({\mu_{t-1}({\bm{x}},\cdot)}).$ (3.3)
Table 4: Values of ${\rm lcb}_{t}({\bm{x}})$ and ${\rm ucb}_{t}({\bm{x}})$ for commonly used robustness measures (Table 3 in Inatsu et al., 2024a ).
@@
$\mathcal{M}(\cdot)$: Monotonic Lipschitz continuous map with Lipschitz constant $K$
@@
Expectation and probability are taken with respect to distribution of ${\bm{w}}$, and $\alpha$-value-at-risk is same meaning as $\alpha$-quantile
@@ 3.3 Acquisition Function
We introduce acquisition functions to determine the next evaluation point $({\bm{x}}_{t},{\bm{w}}_{t})$. In this study, the estimated solution $\hat{\bm{x}}_{t}$ and the next evaluation point ${\bm{x}}_{t}$ are not necessarily identical. In previous studies on BO under robustness considerations and environmental variability (Kirschner et al.,, 2020; Inatsu et al., 2024a, ), ${\bm{x}}_{t}$ is typically chosen based on the upper bound of a credible interval for a robustness measure, while ${\bm{w}}_{t}$ is selected to maximize the posterior variance of $f({\bm{x}}_{t},{\bm{w}})$. We propose a modification to the method for selecting ${\bm{x}}_{t}$ and partially adopt the aforementioned approach. In particular, as ${\bm{w}}_{t}$ is selected based on the posterior variance of $f$, this eliminates the need for hyperparameter tuning in the acquisition function, unlike the credible interval for $F({\bm{x}})$, which depends on a user-specified parameter $\beta_{t}$. To address this, we avoid fixing $\beta_{t}$ explicitly and instead treat it as a realization from a probability distribution. For example, Takeno et al., (2023) proposed IRGP-UCB within the standard BO framework, in which $\beta_{t}$ for GP-UCB is randomly sampled from a two-parameter exponential distribution. Similarly, Inatsu et al., 2024b proposed the randomized straddle method for level-set estimation, where $\beta_{t}$ in the straddle acquisition function (Bryan et al.,, 2005) is drawn from a chi-squared distribution. These methods remove the need to manually specify hyperparameters and also yield tighter theoretical guarantees than conventional GP-UCB or straddle methods. However, these algorithms are designed for problems in which the target function $f$ itself is modeled as a GP, and their theoretical guarantees critically depend on $f$ following GPs. In contrast, the target function in our setting is $F({\bm{x}})$, which generally does not follow a GP, even if $f$ does. Therefore, to derive a theoretically sound acquisition strategy, we modify the GP-UCB method to suit the context in which ${\bm{x}}_{t}$ is selected based on the upper bound of the credible interval for $F({\bm{x}})$. Before introducing this method, we reformulate standard GP-UCB in the conventional BO setting without input uncertainty and highlight its essential structural properties. The following lemma characterizes GP-UCB in this simplified setting:
@@
For a black-box function $f({\bm{x}})$ modeled as a GP, let $\mu_{t-1}({\bm{x}})$ denote the posterior mean, $\sigma^{2}_{t-1}({\bm{x}})$ the posterior variance, and $\beta_{t}\geq 0$ a user-defined parameter. Define ${\rm u}_{t-1}({\bm{x}})=\mu_{t-1}({\bm{x}})+\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}})$ and ${\rm l}_{t-1}({\bm{x}})=\mu_{t-1}({\bm{x}})-\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}})$, and
${\bm{x}}^{({\rm u})}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}{\rm u}_{t-1}({\bm{x}}),\ {\tilde{\bm{x}}^{(f)}_{t}}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm u}_{t-1}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm l}_{t-1}({\bm{x}}))_{+},$
${\hat{\bm{x}}^{(f)}_{t}}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\mu_{t-1}({\bm{x}}),\ {\bm{x}}^{{(f)}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\{{\tilde{\bm{x}}^{(f)}_{t}},{\hat{\bm{x}}^{(f)}_{t}}\}}({\rm u}_{t-1}({\bm{x}})-{\rm l}_{t-1}({\bm{x}}),$
where $(a)_{+}$ denotes $a$ if $a>0$, and otherwise is 0. Then, the equality ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}({\bm{x}}^{{(f)}}_{t}))$ holds.
The proof is provided in Appendix C. Using this lemma, we define the selection rule for ${\bm{x}}_{t}$ as follows.
@@
Let $\xi_{1},\ldots,\xi_{t}$ be independent random variables drawn from the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Assume that ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ are lower and upper bounds that satisfy (3.2), and let $\hat{\bm{x}}_{t}$ be defined by (3.3). Then, ${\bm{x}}_{t}$ is selected as follows:
@@
where $\tilde{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm ucb}_{{t-1}}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{{t-1}}({\bm{x}}))_{+}$.
Here, Lemma 3.1 provides a different perspective on Definition 3.1. The acquisition function in Definition 3.1 appears, at first glance, to be unrelated to the usual GP-UCB. However, the usual GP-UCB can be expressed in the same form as Lemma 3.1, and since this equivalent expression is similar to Definition 3.1, the proposed acquisition function can also be interpreted as an acquisition function based on GP-UCB. Finally, based on ${\bm{x}}_{t}$ selected by (3.4), we determine ${\bm{w}}_{t}$ as follows.
@@
The next environmental variable ${\bm{w}}_{t}$ is selected as follows:
@@ 4 Theoretical Analysis
In this section, we provide theoretical guarantees on the expected regret of the proposed algorithm. Detailed proofs are given in Appendix C. To evaluate the quality of the estimated solution, we define the instantaneous regret $r_{t}$ and cumulative regret $R_{t}$ as follows:
@@
In addition, to derive theoretical guarantees for the proposed method, we introduce the concept of maximum information gain $\gamma_{t}$. This quantity is widely used in the theoretical analysis of GP-based BO and level-set estimation (Srinivas et al.,, 2010; Bogunovic et al.,, 2016; Gotovos et al.,, 2013), and is expressed as follows:
$\gamma_{t}=\frac{1}{2}\sup_{\{({{\bm{x}}_{(1)},{\bm{w}}_{(1)}}),\ldots,({{\bm{x}}_{(t)},{\bm{w}}_{(t)}})\}\subset\mathcal{X}\times\Omega}\log{\rm det}({\bm{I}}_{t}+\sigma^{-2}_{{\rm noise}}\tilde{\bm{K}}_{t}),$
where $({\bm{x}}_{(1)},{\bm{w}}_{(1)}),\ldots,({\bm{x}}_{(t)},{\bm{w}}_{(t)})$ are arbitrary elements of $\mathcal{X}\times\Omega$, $\tilde{\bm{K}}_{t}$ is a $t\times t$ kernel matrix with the $(i,j)$-th entry given by $k({({\bm{x}}_{(i)},{\bm{w}}_{(i)}),({\bm{x}}_{(j)},{\bm{w}}_{(j)})})$. For commonly used kernels, such as the linear, Gaussian, and Matérn kernels, $\gamma_{t}$ is known to grow sublinearly under mild conditions (see, e.g., Theorem 5 in Srinivas et al., (2010)). Let $h(a):[0,\infty)\to[0,\infty)$ be a non-decreasing, concave function satisfying $h(0)=0$, and denote by $\mathcal{H}$ the set of all $h(a)$. We then define a class of functions $q(a):[0,\infty)\to[0,\infty)$, denoted by $\mathcal{Q}$, as follows:
@@
For $\mathcal{Q}$, we impose the following assumption on ${\rm ucb}_{t-1}({\bm{x}}_{t})$, ${\rm lcb}_{t-1}({\bm{x}}_{t})$, and the width term $2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})$.
@@
There exists a function $q(x)\in\mathcal{Q}$ such that for any $t\geq 1,{\bm{x}}_{t},\beta_{t}$, and $\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})$, the following inequality holds:
${\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})\leq q(2\beta^{1/2}_{t}\max_{{\bm{w}}\in\Omega}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})).$ (4.1)
In this study, $q(a)$ plays the same role as $q(a)$ in Inequality (3) of Inatsu et al., 2024a , but while their $q(a)$ does not have any assumptions regarding concave functions, ours does. The reason for this is that, unlike Inatsu et al., 2024a , we are subjecting the expected value of (cumulative) regret to theoretical analysis. This analysis requires an inequality evaluation of the expected value, which they did not need to perform. In particular, we need to use Jensen’s inequality, which requires us to impose additional assumptions regarding concave functions. Similarly to (3.2), it also shares commonalities with Inatsu et al., 2024a in terms of requiring $q(a)$ satisfying (4.1). They provide $q(a)$ for various robustness measures, including the expectation measure. In this study, we impose stronger assumptions on $q(a)$ than those in their; however, since the $q(a)$ provided by Inatsu et al., 2024a satisfies Assumption 4.1 and $q(a)\in\mathcal{Q}$, their results can be applied. Representative robustness measures and their corresponding $q(a)$ functions are summarized in Table 5.444Inatsu et al., 2024a derives $q(a)$ for measures such as the standard deviation and variance, but these rely on constants dependent on the norm in the RKHS, which differ from our setting and are thus excluded. The probabilistic threshold is also omitted as no explicit form of $q(a)$ is provided.
Table 5: Specific forms of $q(a)$ for commonly used robustness measures (modified version of Table 4 in Inatsu et al., 2024a ).
@@
$F({\bm{x}};P)$: Robustness measure $F({\bm{x}})$ defined based on distribution $P$
$q(a;F)$: function $q(a)$ for $F({\bm{x}})$ satisfying Assumption 4.1
$\mathcal{M}(\cdot)$: Monotonic Lipschitz continuous map with Lipschitz constant $K$
$q^{(\cdot)}(a)$: function $q(a)$ for $F^{(\cdot)}({\bm{x}})$ satisfying Assumption 4.1
@@
Expectation and probability are taken with respect to distribution of ${\bm{w}}$,
and $\alpha$-value-at-risk is same meaning as $\alpha$-quantile
@@
Assume that (3.2), the regularity assumption, and Assumption 4.1 hold. Suppose that $\xi_{1},\ldots,\xi_{t}$ are independent random variables following the chi-squared distribution with two degrees of freedom, and that $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Let $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be a function satisfying Assumption 4.1. Then, if Algorithm 1 is performed, the following inequality holds:
@@
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij}}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$. The expectation is taken over all sources of randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
If $\gamma_{t}$ is a sublinear function, the following convergence holds:
@@
Then, if each $h_{i}(a)$ is continuous at 0, it follows from Theorem 4.1 that the $\mathbb{E}[R_{t}]/t$ satisfies
@@
As a special case, according to Table 4 in Inatsu et al., 2024a , $q(a)$ corresponding to the expectation measure satisfies $q(a)=a$; i.e., $n=\zeta_{i}=s_{i}=\lambda_{ij}=\nu_{ij}=1$ and $h_{1}(a)=a$. In this case, $\mathbb{E}[R_{t}]$ satisfies
@@
Similarly, $q(a)$ corresponding to the worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures satisfies $q(a)=2a$ for the mean absolute deviation measure and $q(a)=a$ for the other measures. For these robustness measures, including the expectation measure, the following corollary holds:
@@
Under the assumptions of Theorem 4.1, the following inequality holds for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures:
@@
where $C_{0}=2(2\log(|\mathcal{X}\times\Omega|)+2)/\log(1+\sigma^{-2}_{{\rm noise}})$, and $C$ is 8 for the mean absolute deviation measure and 4 for the other measures.
While Theorem 4.1 and (4.3) provide a guarantee on the expected value of the cumulative regret, they do not directly provide a guarantee on the expected value of the instantaneous regret. Specifically, they do not answer the question regarding which estimated solution $\hat{\bm{x}}_{i}$, for $1\leq i\leq t$, achieves the smallest value of $\mathbb{E}[r_{i}]$. To this end, we define the index $\hat{t}$ of the optimal estimated solution up to time $t$ as follows:
$\hat{t}=\operatornamewithlimits{argmin}_{1\leq i\leq t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})],$ (4.4)
where $\mathbb{E}_{t-1}[\cdot]$ is the conditional expectation given the dataset $D_{t-1}$, defined as follows:
@@
Under the assumptions of Theorem 4.1, the following inequality holds:
@@
where $\hat{t}$ is given by (4.4), and $h_{i}(\cdot)$ along with all coefficients are as defined in Theorem 4.1. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk, and mean absolute deviation measures, the following bound holds:
@@
From Theorem 4.2, if $\gamma_{t}$ is a sublinear function and $h_{i}(\cdot)$ is continuous at 0, then using $h_{i}(0)=0$ and (4.2), $\mathbb{E}[r_{\hat{t}}]$ satisfies
@@
Therefore, the expected regret associated with index $\hat{t}$ converges to zero as $t\to\infty$. However, computing $\hat{t}$ requires solving (4.4), which is generally intractable analytically. Nevertheless, $\hat{t}$ can alternatively be written as follows:
@@
Therefore, $\hat{t}$ can be defined without using the unknown ${\bm{x}}^{\ast}$. The reason for expressing $\hat{t}$ as in (4.4) is to simplify the evaluation of inequalities in theoretical analysis, such as (C.9). Furthermore, the posterior distribution of $f$ given $D_{t-1}$ follows GPs. By using this, we can generate sample paths $\hat{f}_{1},\ldots,\hat{f}_{M}$ from this posterior. For each sample path $\hat{f}_{j}$, we evaluate $F(\bm{x})$ and use these $M$ evaluations to estimate $\mathbb{E}_{t-1}[F({\bm{x}})]$. In this way, $\hat{t}$ can be approximated from the sampled paths. In contrast, for the expectation measure, the following theorem shows that one can use $t$ itself as a substitute for $\hat{t}$.
@@
Under the assumptions of Theorem 4.1, the following holds for the expectation measure: